40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 41     Entry time: Mon Feb 1 12:51:43 2010
Author: Rana, Frank 
Type: DailyProgress 
Category: NoiseBudget 
Subject: First Refcav noise measurement 
Attached is the relative frequency noise between the 2 reference cavities downstairs.

The blue curve shows the calibrated control signal from the PDH loop of the analyzer cavity adjusting the AOM frequency.

The low frequency hump is caused by pointing introduced by the AOM used to lock the second (analyzer) cavity to the first (reference cavity).
Changing the frequency causes causes beam-pointing which, at the moment, is much larger as expected. This might be due to bad alignment of the double-passing beam.

The flat part seems to be the shot noise level of the light to read the analyzer cavity.
We have to increase the power in this part of the setup but have to replace a mirror for that and re-align everything (all EOMs, AOM, cavities etc. That's why we waited until the first result.

Current total power in that part is 1.8mW. The other cavity uses 14.7mW.

The bump at a couple of kHz is due to the PDH loop of the analyzer cavity; phase margin is not enough, but we use one of the unmodified generic PDH boxes so far so it's not a surprise.

Green curve shows the calibrated error point signal. The overall noise performance is ~1000x worse than it should be.
Attachment 1: refcav-noise.pdf  46 kB  | Hide | Hide all
refcav-noise.pdf
ELOG V3.1.3-